OpenStudy (anonymous):

A mortgage broker charges $1000 plus 0.5% of the mortgage amount. What is the mortgage broker fee for an $89,000 mortgage? (3 points)

ganeshie8 (ganeshie8):

so the $1000 is fixed charge

ganeshie8 (ganeshie8):

and 0.5% is the variable charge

ganeshie8 (ganeshie8):

to get the broker fee : find out 0.5% of 89,000 and add it to the fixed charge

OpenStudy (anonymous):

178,000?

OpenStudy (anonymous):

plus the 89,000?

OpenStudy (anonymous):

or no sorry 178,000 + the 1000?

ganeshie8 (ganeshie8):

how did u get 178,000 ?

OpenStudy (anonymous):

I so did this wrong, uh divided 89000 by .5

ganeshie8 (ganeshie8):

think a bit, who on earth wud pay 178,000 to a broker :/

ganeshie8 (ganeshie8):

try again :)

ganeshie8 (ganeshie8):

0.5% of 89,000 = (0.5/100) * 89,000 = ?

OpenStudy (anonymous):

uhhh def no one xD uhm 445?

OpenStudy (anonymous):

so 1445? (:

ganeshie8 (ganeshie8):

445 + 1000 = 1445 is \(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

cool! Thank you! (:

ganeshie8 (ganeshie8):

np :)